OpenStudy (ambuj):

In the 20th row of following triangle 1 2 3 4 5 6 7 8 9 10 ........................... ..........................

OpenStudy (ambuj):

what is its sum?

OpenStudy (jango_in_dtown):

If you look property, consider the 1st column only. the terms are 1,2,4,7 the next term will be 11 and then 16 and so on... See you need to have a formula for a general term of this series

OpenStudy (jango_in_dtown):

*sequence

OpenStudy (jango_in_dtown):

ok, after minor calculations you will find that the general term of the sequence of the 1st column is t_n=(n^2-n+2)/2 so t_20=191

OpenStudy (jango_in_dtown):

now the 20th row will have entries 191,192,...,210

OpenStudy (jango_in_dtown):

so the sum will be 191+192+...+210=(190+1)+(190+2)+(190+3)+...+(190+20)=20*190+(1+2+3+...+20)=20*190+(20)*(20+1)/2=4010

OpenStudy (jango_in_dtown):

@Ambuj